Missing proceeds. Ruth and Betty sell apples at a farmers
market. Ruth’s apples sell at 2 for $1 while Betty’s slightly
smaller apples sell at 3 for $1. When Betty leaves to pick
up her kids they each have 30 apples and Ruth takes
charge of both businesses. To simplify things, Ruth puts all
60 of the apples together and sells them at 5 for $2. When
Betty returns, all of the apples have been sold, but they
begin arguing over how to divide up the proceeds. What
is the problem? Explain what went wrong.

Answer by richwmiller(17219)

30/2=15
30/3=10
10+15=$25
If they each sell their own at their own prices.
60/5=12
12*2=$24
sold in groups of 5 for $2
25-24=1
They are one dollar short.